OpenStudy (anonymous):

Find dimensions of rectangle whose width is 7 miles less than length, and whose area is 78. Please help! I gotta go soon.

OpenStudy (anonymous):

a*(a-7)=78

OpenStudy (anonymous):

Why are you multiplying a?

OpenStudy (anonymous):

Can you give me an equation? I can solve from there, its setting it up that bothers me

OpenStudy (anonymous):

lw=78 l=w-7

OpenStudy (anonymous):

w=l-7****